### v2

# Operational Derivation of Born’s Rule from Causal Consistency in Generalized Probabilistic Theories

Abstract
We present an operational derivation of Born’s rule within finite-dimensional generalized probabilistic theories (GPTs), without presupposing Hilbert-space structure. From a single causal requirement—causal consistency—together with sharp measurements, reversible symmetries, and no-signaling, we show that every state-to-probability map must be affine under mixing; otherwise, non-affine curvature enables superluminal signaling via steering. Affinity plus the standard reconstruction pipeline yields a self-dual quadratic form and forces the probability assignment to coincide with the operational transition $\tau(\psi,\phi)$, which in complex quantum theory is $\tau(\psi,\phi)=|\langle\phi|\psi\rangle|^{2}$.
Our three-stage argument—(i) operational assignment, (ii) causal-consistency constraints, (iii) structural reconstruction—thus recovers complex quantum theory and identifies Born’s rule as a causal fixed point among admissible probabilistic laws. Conceptually, the result shows that the familiar quadratic rule is not a mere artifact of Hilbert-space geometry, but a consequence of demanding that probabilistic predictions remain compatible with relativistic causality in a broad class of GPTs. Limitations: the proof is finite-dimensional and leverages effective self-duality post reconstruction. We also outline steering-based experiments that bound deviations from affinity and render the claim falsifiable in principle.
1 Introduction
The origin of the Born rule $P(i)=|\langle\phi_{i}|\psi\rangle|^{2}$ remains one of the most profound questions in quantum foundations [1, 2]. Unlike classical probability theory, where probabilities often represent subjective ignorance, quantum probabilities appear fundamental and irreducible [15, 16]. This raises the deep question: why does nature employ this specific quadratic probability rule, as opposed to any other map from amplitudes or overlaps to probabilities?
Several derivations exist, but each comes with significant presuppositions. Gleason’s theorem [3] derives the Born rule from noncontextuality but assumes Hilbert-space structure from the outset and requires dimension at least three. Decision-theoretic approaches [4, 5] derive it through rationality arguments but rely on the controversial many-worlds interpretation and axioms that are not purely operational. Zurek’s envariance-based derivation [6] provides physical insights into environment-assisted invariance but still operates fully within the standard quantum formalism, taking the Hilbert-space machinery as given. Relatedly, recent work has explored the status of the Born rule within abstract process-theoretic frameworks. In particular, Agrawal and Wilson [20] showed that any probabilistic process theory can be transformed, via suitable categorical quotient constructions, into an operationally equivalent theory satisfying a generalized Born rule. Their result clarifies the structural relationship between compositional rules and probability assignments in process theories.
Here we take a fundamentally different approach using the framework of Generalized Probabilistic Theories (GPTs) [7, 8, 9, 10]. GPTs provide a broad and flexible mathematical framework that includes classical, quantum, and hypothetical post-quantum theories while maintaining operational clarity [11, 12]. In this picture, states are represented as elements of convex sets, measurements as affine functionals, and composition is constrained by local tomography.
Our key idea is to disentangle two logically distinct steps:
  1. 1.
    the operational assignment of probabilities to outcomes, given some intrinsic notion of overlap or transition probability between states, and
  2. 2.
    the structural reconstruction that identifies the underlying state space with a complex Hilbert space.
Instead of assuming Born’s rule from the outset, we allow for a general functional dependence
$P(\phi|\psi)=\Phi\big(\tau(\psi,\phi)\big),$
where $\tau$ is an operationally defined transition probability and $\Phi$ is a priori an arbitrary admissible function satisfying minimal consistency conditions.
Our key insight is that causal consistency—the impossibility of superluminal signaling—together with an operational notion of affinity under mixing and coarse-graining, uniquely fixes the functional form of quantum probabilities. We prove that in any GPT satisfying six operationally motivated axioms (no signaling, local tomography, purification, continuous reversibility, spectrality, and strong symmetry), and assuming probabilities are affine under mixing/coarse-graining, the only probability assignment compatible with causal consistency is the identity function $P=\tau$ on the operational transition probability. Standard reconstruction theorems then recover complex quantum theory, where $\tau(\psi,\phi)=|\langle\phi|\psi\rangle|^{2}$.
This work makes three main contributions:
  1. 1.
    a causal derivation of the probability rule without presupposing Hilbert space,
  2. 2.
    a clear separation between the operational probability assignment and the recovery of quantum structure, and
  3. 3.
    a demonstration that any non-affine deviation from Born’s rule enables superluminal signaling in the presence of steering.
In this sense, Born’s rule is characterized as a causal fixed point: among all conceivable ways of mapping operational transition probabilities to outcome frequencies, it is the unique one that preserves no-signaling in GPTs with purification and entanglement.
[…]
Figure 1: Logical derivation path from operational principles to Born’s rule. The crucial step is the identification $P=\tau$ from causal consistency and affinity.
2 Operational Framework
In order to state our main result, we briefly recall the standard GPT formalism that we adopt throughout.
Definition 1 (GPT structure [8, 7]).
A physical system is represented by a finite-dimensional real ordered vector space $V$ with a closed generating cone $V^{+}\subset V$ and a distinguished order unit $u\in V^{*}$. We define:
  • States as elements $\omega\in\Omega:=\{\omega\in V^{+}:u(\omega)=1\}$.
  • Effects as affine functionals $e:\Omega\to[0,1]$ with $0\leq e\leq u$, i.e., $e(\omega)\geq 0$ for all $\omega$ and $(u-e)(\omega)\geq 0$ for all $\omega$.
  • Measurements as finite sets $\{e_{i}\}$ of effects with $\sum_{i}e_{i}=u$.
  • Composite systems as locally tomographic: $V_{AB}\simeq V_{A}\otimes V_{B}$, and joint statistics determine the composite state uniquely.
This convex-operational setting encompasses classical (simplex) state spaces, standard quantum theory, and a variety of post-quantum models, and it naturally accommodates entanglement, information-processing tasks, and steering in a unified language [19]. As summarized in Figure 1, the logical structure of our argument proceeds operationally: starting from the GPT axioms, we introduce the transition probability $\tau(\psi,\phi)$, enforce causal consistency and operational affinity, and arrive at the identification $P=\tau$, which in turn enables the reconstruction of complex quantum theory. In particular, extremal states of $\Omega$ represent pure preparations, convex combinations represent classical mixing, and measurements are normalized POVM-like functionals acting on this convex set.
3 Axioms
We adopt six operationally justified axioms, standard in GPT reconstruction programs [9, 10, 8]. For completeness, we restate them here in a compact form.
Axiom 1 (No Superluminal Signaling (NSS) [13]).
For any bipartite experiment on systems $A$ and $B$, local marginals on $A$ are independent of distant measurement choices on $B$:
$P_{A}(a|A,B)=P_{A}(a|A,B^{\prime})$
for all measurements $A$ on system $A$ and $B,B^{\prime}$ on system $B$. An analogous condition holds exchanging $A$ and $B$.
Axiom 2 (Local Tomography (LT) [9]).
Two composite states with identical joint measurement statistics for all local measurements are identical as elements of $V_{A}\otimes V_{B}$. That is, the state of a composite system is completely characterized by the joint outcome probabilities of local measurements on its subsystems.
Axiom 3 (Purification [9]).
Every mixed state $\omega_{A}$ of a system $A$ is the marginal of a pure bipartite state $\Psi_{AB}$, i.e., $\omega_{A}=\mathrm{Tr}_{B}(\Psi_{AB})$, and such a purification is unique up to reversible transformations on $B$.
Axiom 4 (Continuous Reversibility [10]).
The group of reversible transformations on a given system acts transitively and continuously on the set of pure states. In particular, any pure state can be transformed into any other by a reversible transformation along a continuous path.
Axiom 5 (Spectrality / Sharpness [10]).
Every state decomposes into perfectly distinguishable pure states. Moreover, for any such decomposition there exist sharp two-outcome measurements $\{e,e^{\perp}\}$ such that $e(\phi)=1\Rightarrow e^{\perp}(\phi)=0$ for some pure $\phi$, and similarly for the other elements of a maximal distinguishable set.
Axiom 6 (Strong Symmetry [10]).
Any two maximal sets of perfectly distinguishable pure states are connected by a reversible transformation. That is, the theory does not distinguish, at the level of symmetry, between different “classical bases” of a given system.
These six axioms align with established reconstruction programs and suffice to recover the structure of complex quantum theory [9, 10, 8]. They capture the existence of entanglement (via purification), a rich reversible dynamics (continuous reversibility), and a notion of “classical structure” (spectrality and strong symmetry) compatible with information processing and thermodynamics.
4 Operational Transition Probability
We now introduce an intrinsic GPT quantity that generalizes the overlap between pure states, building on operational notions of state discrimination [14].
Definition 2 (Transition probability [10]).
For pure states $\psi,\phi\in\Omega_{\text{pure}}$, define the operational transition probability
$\tau(\psi,\phi):=\sup\{\,e(\psi)\;:\;e(\phi)=1,\;e\text{ is an effect}\,\}.$ (1)
Intuitively, $\tau(\psi,\phi)$ is the maximum probability of mistaking $\psi$ for $\phi$ in an optimal yes/no test designed to accept $\phi$ with certainty. It generalizes the squared inner product $|\langle\phi|\psi\rangle|^{2}$ in standard quantum mechanics but is defined solely in terms of operationally accessible effects. Crucially, this definition is purely operational and makes no reference to any specific probability rule $P$ or to an underlying Hilbert space; it only relies on the convex structure and the existence of sharp measurements.
Lemma 1 (Properties of transition probability).
Under Axioms 3–6, the transition probability $\tau$ satisfies:
  1. 1.
    $0\leq\tau(\psi,\phi)\leq 1$ for all pure states $\psi,\phi$, and $\tau(\psi,\phi)=1$ if and only if $\psi=\phi$.
  2. 2.
    For a sharp two-outcome test $\{e,e^{\perp}\}$ that perfectly distinguishes the pair $\{\phi,\phi^{\perp}\}$, we have
    $e(\psi)=\tau(\psi,\phi),\quad e^{\perp}(\psi)=\tau(\psi,\phi^{\perp}),$
    and hence
    $\tau(\psi,\phi)+\tau(\psi,\phi^{\perp})=1.$
Moreover, once self-duality and bit-symmetry hold (as obtained along the reconstruction path), $\tau(\psi,\phi)$ becomes symmetric: $\tau(\psi,\phi)=\tau(\phi,\psi)$.
Proof sketch.
Property (1) follows immediately from the definition: since $e$ is an effect, we always have $0\leq e(\omega)\leq 1$ for all states $\omega$, hence $0\leq\tau(\psi,\phi)\leq 1$. If $\psi=\phi$, then any effect $e$ with $e(\phi)=1$ yields $e(\psi)=1$, so $\tau(\psi,\phi)=1$. Conversely, if $\tau(\psi,\phi)=1$ with $\psi\neq\phi$, we would obtain an effect that accepts $\psi$ and $\phi$ with probability $1$, contradicting sharpness and the existence of perfectly distinguishing measurements for pure states.
For (2), spectrality and strong symmetry guarantee the existence of sharp two-outcome tests $\{e,e^{\perp}\}$ that distinguish $\phi$ from its perfectly distinguishable partner $\phi^{\perp}$: by definition $e(\phi)=1$ and $e(\phi^{\perp})=0$. Optimality of such tests in the sense of hypothesis testing implies that $e(\psi)$ attains the supremum in the definition of $\tau(\psi,\phi)$, hence $e(\psi)=\tau(\psi,\phi)$. Similarly, $e^{\perp}(\psi)=\tau(\psi,\phi^{\perp})$, and normalization of the two-outcome test gives $e(\psi)+e^{\perp}(\psi)=1$, yielding the claimed additivity relation. Symmetry under exchange of arguments requires self-duality and bit-symmetry, which arise later from the reconstruction theorems and are therefore not assumed at this stage. ∎
5 Main Result: Causal Consistency Fixes the Probability Law
We now formalize the central question: given an operationally defined transition probability $\tau(\psi,\phi)$, how can we assign actual outcome probabilities $P(\phi|\psi)$ in a way that respects causality?
In full generality, we allow that for pure state preparations $\psi$ and pure “target” states $\phi$, the probability of obtaining an outcome associated to $\phi$ is given by
$P(\phi|\psi)=\Phi\big(\tau(\psi,\phi)\big),$ (2)
where $\Phi:[0,1]\to[0,1]$ is a continuous, nondecreasing function satisfying $\Phi(0)=0$ and $\Phi(1)=1$. The standard Born rule corresponds to the special case $\Phi(p)=p$ in quantum theory, but at this stage we do not assume this.
To connect this with the GPT structure and the no-signaling constraint, we impose a natural consistency requirement on $P$ under mixing and coarse-graining.
Operational affinity. For any convex mixture of states $\omega=\lambda\omega_{1}+(1-\lambda)\omega_{2}$ and any effect $e$ (or coarse-grained measurement outcome), operational probability assignments obey
$P(e|\omega)=\lambda P(e|\omega_{1})+(1-\lambda)P(e|\omega_{2}).$ (3)
This expresses the idea that probabilistic mixing at the preparation stage is reflected linearly in the observed outcome frequencies. It is the standard affinity condition used throughout the GPT literature for effects, now extended to the derived map $P$ via $\Phi\circ\tau$.
Theorem 1 (Causal consistency + affinity fix the probability law).
Let $P(\phi|\psi)=\Phi(\tau(\psi,\phi))$ for pure $\psi,\phi$, with $\Phi:[0,1]\to[0,1]$ continuous, nondecreasing, $\Phi(0)=0$, $\Phi(1)=1$. Assume:
  • (i)
    normalization on sharp two-outcome tests, so that for a sharp test distinguishing $\phi$ and $\phi^{\perp}$,
    $P(\phi|\psi)+P(\phi^{\perp}|\psi)=1$
    for all $\psi$;
  • (ii)
    no-signaling (NSS) holds at the level of the full GPT;
  • (iii)
    operational affinity: for any ensemble decomposition and coarse-graining,
    $P(e|\lambda\omega_{1}+(1-\lambda)\omega_{2})=\lambda P(e|\omega_{1})+(1-\lambda)P(e|\omega_{2})$
    for all $\lambda\in[0,1]$.
Then $\Phi(p)=p$ for all $p\in[0,1]$.
Proof sketch.
Normalization on sharp two-outcome tests, together with Lemma 1, implies that for any $p\in[0,1]$,
$\Phi(p)+\Phi(1-p)=1.$
Figure 2 provides a graphical interpretation of this symmetry condition. The figure shows that any admissible deformation of $\Phi$ must satisfy $\Phi(p)+\Phi(1-p)=1$, and highlights how nonlinear deviations—either convex or concave—introduce curvature that later enables superluminal signaling via steering. This visualization clarifies why only the identity map survives the causal-consistency constraint.
Indeed, for a sharp test $\{e,e^{\perp}\}$ distinguishing $\phi$ and $\phi^{\perp}$, we have $\tau(\psi,\phi)=p$ and $\tau(\psi,\phi^{\perp})=1-p$, so
$P(\phi|\psi)=\Phi(p),\quad P(\phi^{\perp}|\psi)=\Phi(1-p),$
and normalization of the two-outcome test yields $\Phi(p)+\Phi(1-p)=1$.
Next, operational affinity under mixing implies that for any pair of pure states $\psi_{1},\psi_{2}$ and $\lambda\in[0,1]$, the probability assigned to a given effect is an affine function of the underlying operational transition probabilities. Using the linear structure induced by convexity and the fact that $\tau$ behaves linearly under appropriate coarse-grainings in sharp measurements, one obtains that $\Phi$ itself must be affine on $[0,1]$, i.e.,
$\Phi(p)=ap+b$
for some real constants $a,b$.
Finally, imposing the boundary conditions $\Phi(0)=0$ and $\Phi(1)=1$ gives $b=0$ and $a=1$, hence $\Phi(p)=p$ for all $p\in[0,1]$. Therefore the only probability law compatible with sharp-test normalization, operational affinity, and no-signaling is the identity map on the operational transition probability, $P(\phi|\psi)=\tau(\psi,\phi)$. ∎
Corollary 1 (Signaling from non-affine $\Phi$).
If $\Phi$ is not affine on $[0,1]$, then there exist two steering-induced ensembles of the same marginal on $B$ and a sharp effect $e$ such that Bob’s local statistics differ, hence no-signaling is violated.
Idea.
By purification (Axiom 3), every mixed state $\omega_{B}$ of subsystem $B$ can be realized as the marginal of a pure bipartite state $\Psi_{AB}$. By steering, different measurement choices on $A$ can generate different ensemble decompositions of the same marginal $\omega_{B}$, say
$\omega_{B}=\sum_{i}p_{i}\psi_{i}=\sum_{j}q_{j}\phi_{j},$
where the ensembles $\{p_{i},\psi_{i}\}$ and $\{q_{j},\phi_{j}\}$ are distinct but yield the same average state on $B$.
Let us fix a sharp effect $e$ on $B$ and define $\tau_{i}=\tau(\psi_{i},\chi)$ and $\sigma_{j}=\tau(\phi_{j},\chi)$ for some reference pure state $\chi$ singled out by $e$ via Lemma 1. Bob’s local statistics under $\Phi$ become
$P_{B}^{(1)}=\sum_{i}p_{i}\Phi(\tau_{i}),\qquad P_{B}^{(2)}=\sum_{j}q_{j}\Phi(\sigma_{j}).$
Since the marginals coincide, we have $\sum_{i}p_{i}\tau_{i}=\sum_{j}q_{j}\sigma_{j}=:\bar{\tau}$. If $\Phi$ is strictly convex or strictly concave on an interval containing $\bar{\tau}$, Jensen’s inequality implies
$\sum_{i}p_{i}\Phi(\tau_{i})\neq\Phi\!\left(\sum_{i}p_{i}\tau_{i}\right)=\Phi(\bar{\tau})=\Phi\!\left(\sum_{j}q_{j}\sigma_{j}\right)\neq\sum_{j}q_{j}\Phi(\sigma_{j}),$
for suitable choices of ensembles. Therefore $P_{B}^{(1)}\neq P_{B}^{(2)}$, meaning that Bob’s local outcome statistics depend on Alice’s ensemble choice, which is in turn controlled by her local measurement setting. This realizes superluminal signaling, contradicting NSS. Hence any non-affine deviation of $\Phi$ leads to operationally detectable signaling. ∎
[…]
Figure 2: Functional constraint from normalization on sharp tests. Combined with operational affinity and boundary conditions, it forces $\Phi(p)=p$. Deviations from affinity correspond to nonlinear curves that generically lead to signaling via steering.
6 Reconstruction to Complex Quantum Theory
Once the identity $P=\tau$ is established, standard reconstruction theorems complete the argument, building on prior work [9, 10, 8]. At this point, the remaining task is to show that GPTs satisfying the six axioms above are equivalent to ordinary complex quantum theory.
Theorem 2 (Reconstruction theorem [9, 10]).
A finite-dimensional GPT satisfying Local Tomography, Purification, Continuous Reversibility, Spectrality, and Strong Symmetry is equivalent to complex quantum theory. In this representation, the operational transition probability takes the form
$\tau(\psi,\phi)=|\langle\phi|\psi\rangle|^{2},$
and the probability rule becomes the standard Born rule.
Proof structure.
We only sketch the main steps, referring to [9, 10, 17, 18] for full details.
  1. 1.
    Self-duality: Purification and the existence of dynamically generated entanglement imply that the state space is self-dual, meaning that there exists an inner product with respect to which the cone of positive states coincides with its dual cone of positive functionals.
  2. 2.
    Spectral decomposition: Spectrality ensures every state has a spectral decomposition into perfectly distinguishable pure states, analogous to the spectral theorem for density operators.
  3. 3.
    Homogeneity and symmetry: Continuous reversibility and strong symmetry imply that the positive cone is homogeneous and that all maximal distinguishable sets are related by reversible transformations. This is the GPT analog of unitary transitivity on orthonormal bases.
  4. 4.
    Classification: Homogeneous self-dual cones endowed with a suitable spectral structure are known to correspond to formally real Jordan algebras [17, 18]. Imposing local tomography excludes real and quaternionic factors, leaving complex Hilbert spaces as the unique consistent option.
  5. 5.
    Transition probability: In complex quantum theory, the operational transition probability between pure states coincides with the squared modulus of the inner product, $\tau(\psi,\phi)=|\langle\phi|\psi\rangle|^{2}$ [10]. Our identification $P(\phi|\psi)=\tau(\psi,\phi)$ then recovers exactly the Born rule.
7 Discussion
7.1 Conceptual Significance
Our derivation reveals that Born’s rule is not merely a mathematical feature of Hilbert spaces, but rather a fundamental consequence of causal consistency in any probabilistic theory satisfying basic operational principles. The identity $P=\tau$ emerges as the unique causal fixed point—the only probability assignment that simultaneously preserves:
  • local causality (no superluminal signaling),
  • linearity under operational mixing, and
  • compatibility with sharp two-outcome measurements on pure states.
In this light, the quadratic nature of quantum probabilities reflects deep constraints imposed by information-processing in a relativistic world, rather than an arbitrary choice of mathematical convenience.
7.2 Comparison with Alternative Derivations
Unlike Gleason’s theorem [3], which assumes Hilbert space structure and noncontextuality from the outset, our derivation begins in the more general GPT framework and recovers the Hilbert space as a consequence rather than a premise. Compared to decision-theoretic approaches [4, 5], our argument does not rely on rationality axioms or the many-worlds interpretation; instead, it rests on causal and operational constraints.
Relative to envariance-based derivations [6], our approach ties the rule directly to information-theoretic and causal requirements, highlighting the role of steering and purification in amplifying deviations from affinity into operational signaling. The guiding principle becomes:
Any nontrivial deformation of Born’s rule generically leaks into the causal structure of the theory.
7.3 Physical Interpretation
The emergence of Born’s rule as a causal fixed point suggests that quantum probabilities reflect robust principles about information processing in a relativistic setting. The specific quadratic form of $P(\phi|\psi)$ appears as the unique law that allows:
  • maximal entanglement and steering,
  • reversible dynamics on pure states,
  • and a rich classical structure encoded in sharp measurements,
without enabling superluminal signaling. From this vantage point, quantum theory is not just a convenient model but the unique causal completion of a broad family of GPTs satisfying natural structural axioms.
7.4 Limitations and Extensions
Our derivation assumes finite dimensionality, excluding continuous-variable systems and field-theoretic generalizations. Extending these results to infinite dimensions would require careful treatment of topology, measure theory, and possibly new axioms compatible with quantum field theory [17]. It also relies on a strong set of assumptions (local tomography, purification, strong symmetry) that might be relaxed in future work.
An important open direction is to identify the minimal subset of axioms needed to force $\Phi(p)=p$. For instance, it would be interesting to understand whether strong symmetry can be weakened, or whether some restricted form of purification suffices. Moreover, one could seek explicit quantitative bounds: how small can deviations from affinity be before they lead to experimentally detectable signaling, given realistic noise and finite statistics?
8 Conclusion
We have established that Born’s rule emerges as a consequence of causal consistency within the GPT framework. Working operationally and without presupposing Hilbert-space structure, we showed that no superluminal signaling, together with normalization on sharp tests and operational affinity under mixing, uniquely fixes the probability assignment to $P(\phi|\psi)=\tau(\psi,\phi)$. Standard reconstruction theorems then recover complex quantum theory, where $\tau(\psi,\phi)=|\langle\phi|\psi\rangle|^{2}$.
Crucially, any non-affine deviation from this rule generically enables superluminal signaling via steering, identifying the Born rule as a causal fixed point among admissible probabilistic laws. In this sense, the universality and robustness of quantum probabilities are not mere artifacts of Hilbert-space geometry, but reflect deep information-theoretic and causal constraints.
Acknowledgments
I thank Jonathan Barrett, Lucien Hardy, and Christopher Fuchs for inspiring work on the GPT framework and foundational questions about probability in quantum theory.